A firm has a profit margin of 6.5% and an equity multiplier of 3.0. Its sales are $440 million, and it has total assets of $220 million. What is its ROE? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Profit margin=net income/sales

Hence net income=440*6.5%=$28.6 million

Equity multiplier=total assets/equity

Hence equity=220/3

=$73.33 million

ROE=net income/equity

=28.6/73.33

=39%